OpenStudy (anonymous):

x+2divided by7-x times x^2-9x+4divided by x^2-5x+6

hartnn (hartnn):

can u factor x^2-5x+6 ?

OpenStudy (anonymous):

\[\frac{ x+2 }{( 7-x )(x^2-9)}/x^2-5x+6\]

OpenStudy (anonymous):

Is that the question?

OpenStudy (anonymous):

no. the x+2 over 7-x times x^2-9x+4 over x^2-5x+6

OpenStudy (anonymous):

1x+1 and 1x-6

hartnn (hartnn):

i think its \(\large \frac{ x+2 }{( 7-x )}\frac{x^2-9x}{x^2-5x+6}\)

OpenStudy (anonymous):

yes^

OpenStudy (anonymous):

there is a +4 after -9x

hartnn (hartnn):

x^2-5x+6 is not (x+1)(x-6), try again :) ohh, so, \(\large \frac{ x+2 }{( 7-x )}\frac{x^2-9x+4}{x^2-5x+6}\)

OpenStudy (anonymous):

how is it not? 1*-6=-6 and 1*1=1 and those added together is -5?

hartnn (hartnn):

instead of +4, is that by any chance, +14 ?

OpenStudy (anonymous):

lol why yes

hartnn (hartnn):

yes, but we need product as +6 now x^2-9x+14 can be factored ^_^

OpenStudy (anonymous):

oh i thought i was looking for the middle, the -5

hartnn (hartnn):

two numbers with product = +6 and sum = -5 ?

OpenStudy (anonymous):

over my head lol

hartnn (hartnn):

how about -2 and -3 ?

OpenStudy (anonymous):

ok =]

OpenStudy (anonymous):

x-2, x-3

OpenStudy (anonymous):

top is -7, -2?

hartnn (hartnn):

correct! so you can factor the quadratic equations :) now try to factor, x^2-9x+14 :)

hartnn (hartnn):

yes, -7, -2 are the 2 numbers, so factored form will be ?

OpenStudy (anonymous):

(x-7)(x-2)

OpenStudy (anonymous):

is that what your asking?

hartnn (hartnn):

yes, correct now plug these factored forms in to your original expression and see what gets cancelled ! note : 7-x = - (x-7)

OpenStudy (anonymous):

is it (x+2)/(x+3)

hartnn (hartnn):

the final answer ? no... i think only x-7 is getting cancelled, everything else remains...

OpenStudy (anonymous):

the x-2's dont cancle?

hartnn (hartnn):

\(\large \frac{ x+2 }{-(x-7 )}\frac{(x-2)(x-7)}{(x-2)(x-3)}\) and yes, x-2's also cancel

hartnn (hartnn):

\(\large \frac{ (x+2) }{-\cancel{(x-7 )}}\frac{\cancel{(x-2)}\cancel{(x-7)}}{\cancel{(x-2)}(x-3)}\) what remains ?

OpenStudy (anonymous):

thank you!

hartnn (hartnn):

finally we have -(x+2)/(x-3) or (x+2)/(3-x) welcome ^_^